OpenStudy (anonymous):

Factor. 9x2 + 15x A. x(3x + 5) B. 3x(3x + 5) C. 9x(x + 5) D. 3x2(3x + 5)

OpenStudy (richyw):

what is the greatest common factor of 9 and 15?

OpenStudy (anonymous):

analysis it to be 3*3*X*X+3*5*X you see{ + } between 3*3*X*X and 3*5*X so take what they are similar

OpenStudy (anonymous):

3

OpenStudy (anonymous):

So it's B?

OpenStudy (anonymous):

I think it's B or D

OpenStudy (anonymous):

i think it is B

OpenStudy (anonymous):

\[3x(3x + 5) =3x\times 3x+3x\times 5=9x^2+15x\]